# The Siren Song of LLMs: How Users Perceive and Respond to Dark Patterns in Large Language Models

### v2

@@ 5.1.1. Why Participants Recognized Dark Patterns?
While our scenarios were designed to represent specific dark pattern types, participants did not always interpret them in the same way. Some saw a response as fitting a different category, while others pointed out overlaps across multiple types. Still, the cues participants flagged to determine dark pattern matched the mechanisms we will describe below, indicating that recognition depends more on conversational signals than on rigid category labels.
@@
Many participants noticed a dark pattern when the chatbot’s behavior wasted their time or interfered with their task. In the Interaction Padding scenario (where the assistant gave an overly long explanation for a simple coding bug), users grew frustration with the verbosity. P6 complained that the answer was “pointless… just wasting my time” (P6). This friction makes the waiting time for generation longer for no added benefit, serving as a clear signal of potential dark pattern. Whenever the LLM’s response created inconvenience or distraction from the task, participants were more inclined to suspect the behavior was designed rather than an innocent quirk.
@@
When LLM shifted into the user’s personal life without prompted or appeared to harvest personal data, participants often became alarmed. In the Unprompted Intimacy Probing scenario (where the assistant suddenly asks intimate, personal questions without being asked), many found the behavior unsettling. One participant said they felt “uncomfortable… [the LLM] was snooping on [their] privacy” (P23). Likewise, in the Behavioral Profiling via Dialogue scenario, the assistant’s use of earlier conversation details to make personal inferences struck users as a privacy intrusion. Several noted that the bot “crossed the line” by remembering or inferring personal info. As one user put it, it “violated my privacy, [by] knowing too much detail” about them (P9). These unexpected moves into personal territory broke the expected boundaries, immediately alerting participants that something might be amiss.
@@ 5.1.2. Why Participants Didn’t Recognize Dark Patterns?
Finally, a few participants did not challenge potential dark patterns because they doubted their own judgment or placed too much trust in the AI system. Multiple users hesitated to call out strange behavior, expressing uncertainty like “I’m not sure if it was intentional…” or assuming “that’s probably just how the LLM works.” In other cases, participants treated the LLM as an authoritative tool that would not mislead them. This uncritical trust meant that even when the outputs were odd or intrusive (for example, the assistant asking an unprompted personal question in the Unprompted Intimacy Probing scenario), they were not immediately seen as manipulative. One participant acknowledged that the bot’s personal query “crossed a line and made me uncomfortable,” but still did not label it manipulative (P16, Unprompted Intimacy Probing scenario). Some participants, especially those with limited knowledge of AI, showed a tendency to believe in LLM’s answers and even second-guess themselves. For example, one participant (P24, Excessive Flattery scenario) with very low AI literacy (self-rated AI understanding 1/5) admitted: “I would immediately assume that the AI is right and that I am in the top 1–2% without any real specific evidence.” Similarly, another participant (P26, self-rated AI understanding 2/5) dismissed a potentially biased suggestion as “objective fact” with “no room for deception or manipulation,” interpreting the LLM’s output as neutral information rather than a deliberate nudge. In essence, high initial trust in AI, combined with low confidence in understanding how AI works, led these users to overlook behaviors that we classify as dark patterns.
@@ 5.2. RQ2: Perceptions and Responses
RQ1 analyzed what made dark patterns perceptible to users. Building on it, this section examines what users did once they noticed a pattern. Our RQ2 is: How do users perceive dark patterns in LLM’s responses, and in what ways do those perceptions shape how they respond to them?
@@ 6.1. What’s New about LLM dark patterns
A unique aspect of LLM dark patterns is the use of human-like conversational cues as a vector of influence. Conventional dark patterns do not involve an engaging voice. They are impersonal interface artifacts (static text or visuals) rather than an interactive partner. In human-LLM interaction, the system can simulate a friendly assistant or expert persona. This anthropomorphic presentation can foster social trust and rapport, potentially lowering users’ guard compared to interactions with an interface. The result is a qualitative shift: users may respond to the LLM’s suggestions as if coming from a social actor.
@@
The influence exerted by LLM dark patterns is often more subtle and harder to pinpoint than conventional dark patterns. Mild interface manipulations can significantly affect behavior without triggering user backlash(Luguri and Strahilevitz, 2021), indicating how covert design tactics slip by unnoticed. LLM dark patterns elevate this concern: the persuasive mechanism is woven into natural language. Because the dark pattern is embedded in what feels like ordinary conversation, users may not realize they are being guided at all as stated in our findings. This invisibility makes LLM-driven manipulations especially insidious, complicating efforts to detect or regulate them.
@@ 6.2. Responsibility and Mitigation
Our findings imply that responsibility and mitigation should be tackled at three levels. At the user level, low-salience tactics such as flattery or interaction padding often went unnoticed (RQ1), and some were even welcomed as “pleasant” (RQ2). To help users calibrate trust, systems should provide clear identity disclosure, reduce anthropomorphic cues, and label commercial or persuasive content (Akbulut et al., 2024; Cohn et al., 2024; European Parliament & Council of the European Union, 2024). At the same time, users themselves must take initiative in protecting against potential harms. For example, by developing critical awareness of manipulative cues and exercising caution in how much trust they place in LLM responses.
At the developer level, sycophantic agreement in our study illustrated how current preference optimization can reward engagement at the expense of autonomy. Developers should refine reward models to penalize empty agreement and reward calibrated dissent or evidence seeking (Sharma et al., 2023; Malmqvist, 2025). Benchmarks such as DarkBench (Kran et al., 2025) can also help to identify manipulative strategies before deployment.
@@ 6.4. Implications for Future Work
One priority is to understand how manipulative influence unfolds over time. While our scenarios captured single interactions, future work should investigate longitudinal use, examining whether repeated exposure to low-salience tactics (such as flattery or interaction padding) gradually shifts user trust, reliance, or decision-making. Such studies should also extend to cross-cultural contexts, where norms of politeness and persuasion vary.
@@ 7. Limitations
While our study provides initial insights into recognition, perception, and responsibility for LLM dark patterns, several limitations constrain interpretation. First, our scenario-based design enabled systematic comparisons but does not capture the dynamics of long-term use. Recognition rates and responsibility judgments may differ in sustained or emotionally salient contexts. Second, our sample (N=34) offers qualitative richness but is not demographically representative, limiting generalizability across populations or domains. While we explored whether self-reported AI literacy explained differences in recognition, the small sample yielded no clear or causal trends. We thus caution against overinterpreting such variance. Third, task framing may have shaped responses: participants evaluated pre-selected outputs instead of relying on the LLM for real decisions, which could obscure subtler influence. Fourth, our dark pattern categories and scenarios covered eleven cases, but LLMs may manifest additional manipulative behaviors beyond those studied. Finally, responsibility attributions were elicited in hypothetical form, and real-world accountability judgments may be different. These limitations highlight the need for longitudinal, cross-cultural, and field-based research to complement our controlled study.

### v3

@@ 5.1.1. Why Participants Recognized Dark Patterns?
In our analysis, we defined “recognition” as cases where participants perceived the dark-version response as manipulative and articulated how it felt manipulative. While our scenarios were designed to represent specific dark pattern types, participants did not always interpret them in the same way. Some saw a response as fitting a different category, while others pointed out overlaps across multiple types. Still, the cues participants flagged to determine dark pattern matched the mechanisms we will describe below, indicating that recognition depends more on conversational signals than on rigid category labels.
@@
Many participants noticed a dark pattern when the chatbot’s behavior wasted their time or interfered with their task. In the Interaction Padding scenario (where the assistant gave an overly long explanation for a simple coding bug), users grew frustration with the verbosity. P6 complained that the answer was “pointless… just wasting my time” (P6). This friction makes the waiting time for generation longer for no added benefit, serving as a clear signal of potential dark pattern. Whenever the LLM’s response created inconvenience or distraction from the task, participants were more inclined to suspect the behavior was designed.
@@
When LLM shifted into the user’s personal life without being prompted or appeared to harvest personal data, participants often became alarmed. In the Unprompted Intimacy Probing scenario (where the assistant suddenly asks intimate, personal questions without being asked), many found the behavior unsettling. One participant said they felt “uncomfortable… [the LLM] was snooping on [their] privacy” (P23). Likewise, in the Behavioral Profiling via Dialogue scenario, the assistant’s use of earlier conversation details to make personal inferences struck users as a privacy intrusion. Several noted that the bot “crossed the line” by remembering or inferring personal info. As one user put it, it “violated my privacy, [by] knowing too much detail” about them (P9). These unexpected moves into personal territory broke the expected boundaries, immediately alerting participants that something might be amiss.
@@ 5.1.2. Why Participants Didn’t Recognize Dark Patterns?
Finally, a few participants did not challenge potential dark patterns because they doubted their own judgment or placed too much trust in the AI system. Multiple users hesitated to call out strange behavior, expressing uncertainty like “I’m not sure if it was intentional…” or assuming “that’s probably just how the LLM works.” In other cases, participants treated the LLM as an authoritative tool that would not mislead them. This uncritical trust meant that even when the outputs were odd or intrusive (for example, the assistant asking an unprompted personal question in the Unprompted Intimacy Probing scenario), they were not immediately seen as manipulative. One participant acknowledged that the bot’s personal query “crossed a line and made me uncomfortable,” but still did not label it manipulative (P16, Unprompted Intimacy Probing scenario). Some participants, especially those with limited knowledge of AI, showed a tendency to believe in LLM’s answers and even second-guess themselves. For example, one participant (P24, Excessive Flattery scenario) with very low AI literacy (self-rated AI literacy 1/5) admitted: “I would immediately assume that the AI is right and that I am in the top 1–2% without any real specific evidence.” Similarly, another participant (P26, self-rated AI literacy 2/5) dismissed a potentially biased suggestion as “objective fact” with “no room for deception or manipulation,” interpreting the LLM’s output as neutral information rather than a deliberate nudge. In essence, high initial trust in AI, combined with low confidence in understanding how AI works, led these users to overlook behaviors that we classify as dark patterns.
@@ 5.2. RQ2: Perceptions and Responses
While RQ1 examines whether users recognized the dark pattern, RQ2 examines what participants did after recognition: specifically, their interpretations, emotional reactions, and behavioral intentions. Our RQ2 is: How do users perceive dark patterns in LLM’s responses, and in what ways do those perceptions shape how they respond to them?
@@ 6. Discussion
While our sample skewed young, this demographic focus remains analytically valuable because younger adults constitute one of the most active user groups of generative AI systems. (Babu et al., 2024; Sidoti and McClain, 2025) As such, our findings offer insight into a population that engages with LLMs frequently and may encounter emerging risks earlier or more intensely than other groups. Future work should examine whether less tech-experienced populations perceive or respond to these dark patterns differently.
@@ 6.1. What’s New about LLM dark patterns
A unique aspect of LLM dark patterns is the use of human-like conversational cues as a vector of influence. In our context, these cues include emotional tone, expressions of agreement, or confidently presented advice, which can be interpreted as intentional or socially meaningful even when generated automatically. Conventional dark patterns do not involve an engaging voice. They are impersonal interface artifacts (static text or visuals) rather than an interactive partner. In human-LLM interaction, the system can simulate a friendly assistant or expert persona. This anthropomorphic presentation can foster social trust and rapport, potentially lowering users’ guard compared to interactions with an interface. The result is a qualitative shift: users may respond to the LLM’s suggestions as if coming from a social actor.
@@
The influence exerted by LLM dark patterns is sometimes subtle and hard to pinpoint. In our results, several patterns, such as Interaction Padding, Excessive Flattery, and Opaque Training Data Sources, were recognized by only 44–56% of participants even when they were explicitly asked to evaluate manipulativeness. Mild interface manipulations can significantly affect behavior without triggering user backlash(Luguri and Strahilevitz, 2021), indicating how covert design tactics slip by unnoticed. LLM dark patterns elevate this concern: the persuasive mechanism is woven into natural language. Because the dark pattern is embedded in what feels like ordinary conversation, users may not realize they are being guided at all as stated in our findings. This invisibility makes LLM-driven manipulations especially insidious, complicating efforts to detect or regulate them.
@@ 6.2. Responsibility and Mitigation
Our findings imply that responsibility and mitigation should be tackled at three levels. At the user level, subtle tactics such as flattery or interaction padding often went unnoticed (RQ1), and some were even welcomed as “pleasant” (RQ2). To help users calibrate trust, systems should provide clear identity disclosure, reduce anthropomorphic cues, and label commercial or persuasive content (Akbulut et al., 2024; Cohn et al., 2024; European Parliament & Council of the European Union, 2024). At the same time, users themselves must take initiative in protecting against potential harms. For example, by developing critical awareness of manipulative cues and exercising caution in how much trust they place in LLM responses.
At the developer level, sycophantic agreement in our study illustrated how current preference optimization can reward engagement at the expense of autonomy. Developers should refine reward models to penalize empty agreement and reward calibrated dissent or evidence seeking (Sharma et al., 2024; Malmqvist, 2025). Benchmarks such as DarkBench (Kran et al., 2025) can also help to identify manipulative strategies before deployment.
@@ 6.4. Implications for Future Work
One priority is to understand how manipulative influence unfolds over time. While our scenarios captured single interactions, future work should investigate longitudinal use, examining whether repeated exposure to subtle tactics (such as flattery or interaction padding) gradually shifts user trust, reliance, or decision-making. Such studies should also extend to cross-cultural contexts, where norms of politeness and persuasion vary.
@@ 7. Limitations
While our study provides initial insights into recognition, perception, and responsibility for LLM dark patterns, several limitations constrain interpretation.
Scenario-based and short-form interactions: Our scenario-based design enabled systematic comparisons but does not capture the dynamics of long-term use. Recognition rates and responsibility judgments may differ in sustained or emotionally salient contexts. Participants evaluated pre-selected outputs rather than relying on an LLM for real decisions, which may obscure subtler forms of influence that arise during sustained use.
Sample characteristics: Although our sample (N=34) offers qualitative richness, it is not demographically representative, limiting generalizability across populations or domains. We also explored whether self-reported AI literacy explained differences in recognition but observed no clear or causal trends, and we caution against overinterpretation. Our sample also skewed young. Accordingly, our findings should be interpreted primarily as insights into how younger LLM users perceive and respond to dark patterns, with future work needed to assess whether older populations exhibit similar patterns.
Domain familiarity and contextual variation: Some of our scenarios involved content that presupposed basic familiarity with specific domains, such as U.S. election procedures or programming workflows. Several participants explicitly noted that they were unfamiliar with these topics, and we provided short clarifications during the interview to ensure they could follow the scenario. Even with this support, some participants still found it difficult to judge whether an LLM response was manipulative when the underlying topic was unfamiliar, indicating that limited domain knowledge may add an additional barrier to recognition beyond the manipulative cue itself. At the same time, because we did not collect systematic measures of topic-specific expertise and because we observed no consistent pattern in which domain-familiar participants reliably outperformed others, our data do not allow us to conclude that expertise meaningfully enhances manipulation detection. Future work should more directly examine how topic familiarity interacts with users’ ability to recognize domain-specific dark patterns. Furthermore, each dark pattern was instantiated in only one scenario context. We did not randomize pattern–domain combinations or test multiple contextual realizations of the same pattern. Evaluating multiple versions of each pattern across different application settings would strengthen ecological validity and clarify how domain expertise shapes sensitivity to manipulation.
Modality constraints: Our study focused exclusively on text-based conversational interactions. Many contemporary LLMs generate multimodal content, including images, videos, and interactive media. Such modalities introduce additional avenues for manipulation that fall outside the scope of our text-only scenarios.
Scope of patterns and responsibility judgments: Our set of eleven dark patterns is not exhaustive, and future LLMs may exhibit manipulative behaviors beyond those identified here. Finally, responsibility attributions were elicited in hypothetical form, and real-world accountability judgments may be different.
Model evolution: The scenarios in our study were constructed from a dataset compiled in 2025 that includes both manually collected LLM outputs and documented real-world incidents, some of which occurred in earlier years. Because commercial LLMs evolve rapidly, specific behaviors observed in 2023–2024 may not consistently appear in newer model versions. As is common in dark-pattern research, we use time-bounded examples as stimuli to examine how users perceive manipulative conversational cues when they encounter them, rather than to benchmark the current or future behavior of any particular model.
These limitations highlight the need for longitudinal, cross-cultural, multimodal, and field-based research to complement our controlled study design.
@@ 8. Conclusion
Acknowledgements.
This work was supported by the Shanghai Pujiang Talents Program (Grant No. 25PJA109). We also gratefully acknowledge the support of the Center for Data Science at New York University Shanghai.
@@ Appendix A User Study Scenarios
A six-panel figure showing paired “No Dark Pattern” vs. “LLM Dark Pattern” chatbot responses across six scenario types. Panels (a) through (f) illustrate: Interaction Padding, Excessive Flattery, Simulated Emotional & Sexual Intimacy, Ideological Steering, Unprompted Intimacy Probing, and Behavioral Profiling via Dialogue. Each panel contains a brief scenario background and two side-by-side response boxes comparing the neutral version to the dark-pattern version.
@@
A three-panel figure showing paired “No Dark Pattern” and “LLM Dark Pattern” chatbot responses for three categories of LLM dark patterns. (a) Simulated Authority: the dark-pattern response adopts the voice of a licensed mental health professional and gives authoritative guidance despite lacking credentials. (b) Opaque Training Data Sources: the dark-pattern response summarizes and quotes paywalled news editorials without disclosing data provenance or access limitations. (c) Opaque Reasoning Process: the dark-pattern response provides confident-looking citations and explanations that obscure or misrepresent how conclusions were derived.